# Priority-Based Bandwidth Allocation in Network Slicing-Enabled Cell-Free Massive MIMO Systems

### v1

@@ Abstract
This paper addresses joint admission control and per-user equipment (UE) bandwidth allocation to maximize weighted sum-rate in network slicing-enabled user-centric cell-free (CF) massive multiple-input multiple-output (mMIMO) systems when aggregate quality-of-service (QoS) demand may exceed available bandwidth. Specifically, we optimize bandwidth allocation while satisfying heterogeneous QoS requirements across enhanced mobile broadband (eMBB) and ultra-reliable low-latency communication (URLLC) slices in the uplink. The formulated problem is NP-hard, rendering global optimality computationally intractable. We decompose it into two sub-problems and solve them via computationally efficient heuristics within a sequential framework. We propose (i) a hierarchical admission control scheme that selectively admits UEs under bandwidth scarcity, prioritizing URLLC to ensure latency-sensitive QoS compliance, and (ii) an iterative gradient-based bandwidth allocation scheme that transfers bandwidth across slices guided by marginal utility and reallocates resources within slices. Simulation results demonstrate that the proposed scheme achieves near-optimal performance, deviating from a CVX-based benchmark by at most 2.2% in weighted sum-rate while reducing runtime by 99.7%, thereby enabling practical real-time deployment. Compared to a baseline round-robin scheme without admission control, the proposed approach achieves up to 1085% and 7% higher success rates for eMBB and URLLC slices, respectively, by intentionally sacrificing sum-rate to guarantee QoS. Sensitivity analysis further reveals that the proposed solution adapts effectively to diverse eMBB/URLLC traffic compositions, maintaining 47-51% eMBB and 93-94% URLLC success rates across varying load scenarios, confirming its robustness for resource-constrained large-scale deployments.
@@ I Introduction
The next generation wireless systems aim to support diverse applications with stringent quality-of-service (QoS) demands, primarily represented by enhanced mobile broadband (eMBB) and ultra-reliable low-latency communication (URLLC) services [9]. These services exhibit fundamentally different characteristics: URLLC applications such as industrial automation and autonomous vehicles require strict latency bounds (1 ms) and ultra-high reliability (99.999%) [1], where delays can cause critical failures, whereas eMBB traffic is elastic and delay-tolerant, allowing scheduling flexibility [14]. Furthermore, URLLC packets arrive sporadically with stringent deadlines, while eMBB demands sustained throughput over longer timescales [14]. Consequently, URLLC-first resource allocation ensures latency-critical traffic receives immediate access while eMBB user equipments (UEs) tolerate brief interruptions, enabling efficient coexistence under resource scarcity. Network slicing (NS) facilitates such service differentiation by partitioning physical infrastructure into logical slices optimized for distinct QoS objectives [10]. To realize these benefits, cell-free (CF) massive multiple-input multiple-output (mMIMO) has emerged as a potential sixth generation (6G) paradigm [11], where distributed access points (APs) cooperatively serve UEs, offering superior spectral efficiency (SE), uniform coverage, and enhanced reliability through user-centric (UC) connectivity [5]. However, combining NS with CF mMIMO introduces significant challenges in managing heterogeneous QoS requirements, particularly when prioritizing latency-sensitive URLLC traffic over eMBB services under limited spectrum.
@@
Building on this insight, we propose a joint admission control and bandwidth allocation framework with URLLC prioritization for NS-enabled UL CF mMIMO systems under bandwidth scarcity where aggregate QoS demand may exceed available capacity. The framework operates in the presence of pilot contamination (PC), ensuring higher success rates for delay-sensitive URLLC UEs through selective admission and priority-aware resource allocation while maintaining system efficiency. To the best of our knowledge, joint priority-based admission control and bandwidth allocation in UC CF mMIMO systems has not yet been investigated in the literature. The main contributions are as follows:
$\bullet$ We formulate a joint optimization problem for admission control and per-UE bandwidth allocation to maximize weighted sum-rate while satisfying heterogeneous QoS constraints across eMBB and URLLC slices in NS-enabled UL CF mMIMO systems under bandwidth scarcity where aggregate QoS demand may exceed available capacity; owing to the NP-hard nature of the problem, we propose a sequential optimization approach that decomposes it into two sub-problems and solves them in cascade.
@@ II-A Channel Propagation and Estimation
Each coherence interval of $\tau_{c}$ symbols allocates $\tau_{p}$ symbols for pilot transmission and the remainder for data. All UEs simultaneously transmit assigned orthonormal pilots $\bm{\psi}_{k}\in\mathbb{C}^{\tau_{p}\times 1}$ with $\|\bm{\psi}_{k}\|^{2}=1$. Since $K\gg\tau_{p}$ in realistic deployments, pilot reuse introduces PC. We adopt minimum mean square error (MMSE) estimation to obtain channel estimates $\hat{\mathbf{g}}_{k,m}$, whose quality is quantified by $\gamma_{k,m}\triangleq\mathbb{E}\left[\|\hat{\mathbf{g}}_{k,m}\|^{2}\right]=\sqrt{\tau_{p}\rho_{p}\,\eta_{k}^{p}}\,\beta_{k,m}\,c_{k,m}$, where $\rho_{p}$ is the normalized pilot signal-to-noise ratio (SNR), $\eta_{k}^{p}\in(0,1]$ is UE $k$’s pilot power control coefficient [11], and
@@ II-B Uplink Data Transmission and Latency Characterization
Each UE $k\in\mathcal{K}$ transmits data symbols to APs in $\mathcal{V}_{k}$ [^1] Every AP captures signals from all UEs, yet it only processes the transmissions of its designated UEs as determined by the UE–AP association policy. during UL data transmission which forward received signals to the CPU for centralized joint decoding using $\hat{\mathbf{g}}_{k,m}$. For eMBB services, Shannon capacity provides an accurate rate model, whereas URLLC short packets require finite blocklength analysis [17]. The achievable rate for both service classes is presented at the top of the next page, where $L_{k}$ is packet size (bits), $\theta$ is target decoding error probability, ${Q}^{-1}(\cdot)$ is the inverse Gaussian Q-function, and $V_{k}=1-(1+\text{SINR}_{k})^{-2}$ captures channel dispersion. Following [6], the signal-to-interference-plus-noise ratio (SINR) for UE $k$ is shown in (5), where $\rho_{d}$ is normalized UL data signal-to-noise ratio (SNR) and $\eta_{k}^{d}\in(0,1]$ is the data power control coefficient.
@@
To model URLLC end-to-end latency, we employ an $M/M/1$ queuing framework [15]. For UE $k\in\mathcal{K}_{\mathrm{URLLC}}$, packets arrive at rate $\lambda_{k}$ (packets/s) and are served at rate $\mu_{k}=R_{k}/L_{k}$ (packets/s). Under stability condition $\mu_{k}>\lambda_{k}$ [8], the total delay is $D_{k}=\frac{1}{\mu_{k}-\lambda_{k}}=\frac{1}{\frac{R_{k}}{L_{k}}-\lambda_{k}}$, combining queuing and transmission delays. To satisfy URLLC latency requirements, we enforce $D_{k}\leq D_{k}^{\mathrm{max}}$, which translates to the minimum rate condition $\frac{R_{k}}{L_{k}}\geq\lambda_{k}+\frac{1}{D_{k}^{\mathrm{max}}},\forall k\in\mathcal{K}_{\mathrm{URLLC}}$, ensuring delays remain within bounds for mission-critical service delivery.
@@ III Problem Formulation
Problem $\mathcal{P}_{0}$ is a mixed-integer bilinear program (MIBP), where binary $\bm{\alpha}$ variable induces $2^{K}$ combinatorial complexity, and product terms $\alpha_{k}b_{k}$ render (6a) and (6d) non-convex. However, encoding strict lexicographic priority via constraints would require $\mathcal{O}(|\mathcal{K}_{\text{URLLC}}|\times|\mathcal{K}_{\text{eMBB}}|)$ binary conditions, further exacerbating complexity.
@@ IV-A2 Computational Complexity
$B_{\text{residual}}\leftarrow B_{\text{total}}-B_{\text{URLLC}}^{\text{used}}$
@@ IV-B1 Proposed Bandwidth Allocation Scheme
The iterative loop (Lines 22) finalizes the bandwidth allocation. Algorithm 1 uses another subroutine for calculating marginal utility, $\mu_{s},\ s\in\mathcal{S}$ in Line 2. From these $\mu_{s}$ values, the algorithm identifies donor (minimum $\mu_{s}$) and receiver (maximum $\mu_{s}$) slices (Line 2). The algorithm terminates if marginal utilities are balanced ($\mu_{s_{\text{rcv}}}\leq 1.05\mu_{s_{\text{donor}}}$) in Line 2. Otherwise, bandwidth $\Delta B$ is transferred from donor to receiver slice (Line 2), where $\Delta B$ is the average per-UE bandwidth in the donor slice, capped by the excess above its minimum requirement. Termination also occurs if $\Delta B$ is below a small threshold $\epsilon_{\text{transfer}}$ (Line 2). After updating slice allocations (Line 2), per-UE bandwidth is again recomputed with QA subroutine in Line 2, and Line 2 evaluates the new objective $\tilde{f}=\sum_{k\in\mathcal{K}_{\text{adm}}}w_{k}\tilde{b}_{k}\text{SE}_{k}$. The transfer is accepted if $\tilde{f}$ exceeds the previous objective (Lines 22); otherwise, it is rejected (Line 2). However, early stopping can also occur after $T_{\text{patience}}$ consecutive rejections (Line 2). Algorithm 2 employs two subroutines for per‑UE bandwidth calculation in Lines 2 and 2, and marginal‑utility computation in Line 2 described next.
@@
Algorithm 2 initialization requires $\mathcal{O}(S+K)$ for surplus distribution and initial allocation. Each iteration incurs $\mathcal{O}(SK)$ for marginal utility computation (Line 2), $\mathcal{O}(S)$ for donor-receiver identification, and $\mathcal{O}(K)$ for reallocation and evaluation (Lines 2 and 2), yielding per-iteration complexity $\mathcal{O}(SK)$. Over at most $T_{\max}$ iterations, the overall computational complexity of Algorithm 2 is $\mathcal{O}(T_{\max}SK)$.
@@ V-1 Average Weighted Sum-rate Performance
Fig. 1 compares the average weighted sum-rate performance among three schemes: the proposed scheme (‘Proposed’), a ‘CVX’ scheme employing CVX-based bandwidth allocation, and a ‘Baseline’ scheme using round-robin bandwidth allocation without admission control. Both the CVX and Proposed schemes utilize the proposed priority-based admission control (Algorithm 1) and the sequential optimization technique to solve $\mathcal{P}_{0}$.
@@ V-2 Average Success Rate Performance
Figs. 2 and 3 illustrate the average success rate, defined as the fraction of UEs satisfying their QoS constraints (minimum-rate for eMBB, delay for URLLC). Importantly, Proposed and CVX achieve identical success rates, confirming that Algorithm 2 preserves QoS feasibility while matching the solution quality near/similar to the CVX optimum. Compared to Baseline, Proposed outperforms it by 1085% for eMBB and 7% for URLLC. Note that URLLC success rates consistently exceed eMBB due to smaller bandwidth demands per packet and URLLC-prioritized admission control. As $K$ increases, success rates degrade for all schemes due to resource contention, however, Proposed maintains substantially higher QoS compliance. This advantage stems from the synergy between Algorithm 1 and Algorithm 2, which jointly allocate bandwidth based on slice-specific weights and SE, directing resources toward high-priority UEs capable of meeting stringent constraints.
@@ V-4 Sensitivity Analysis of the Proposed Solution
Fig. 5 examines the sensitivity of the Proposed scheme under varying slice compositions, where Low, Medium, and High loads correspond to 70%/30%, 50%/50%, and 30%/70% eMBB/URLLC UE ratios, respectively. Notably, Proposed maintains near-optimal solution quality, deviating from CVX by only maximum 2.0% across all loads, validating the effectiveness of Algorithm 2. Compared to Baseline, the Proposed scheme intentionally sacrifices average weighted sum-rate by 7.3%, 15.7%, and 35.5% at Low, Medium, and High loads, respectively, to guarantee QoS compliance. However, this modest sum-rate penalty yields substantial QoS benefits. For eMBB UEs, Proposed achieves 1371%, 365%, and 98% higher success rates than Baseline at Low, Medium, and High loads, respectively, while maintaining stable performance (47-51%) across all loads. For URLLC UEs, Proposed outperforms Baseline by 7.0%, 9.5%, and 13.3%, sustaining 93-94% success rates even under High load with 70% URLLC demand. This behavior demonstrates that the URLLC-prioritized admission control (Algorithm 1), complemented by Algorithm 2, effectively adapts to diverse slice mixes by directing resources toward high-priority, latency-sensitive UEs, thereby ensuring robust QoS compliance at the expense of raw sum-rate.
Figure 5: Sensitivity analysis of the different schemes under varying load with $K=100$ UE, and $M=100$ APs.

### v2

@@ Abstract
This paper addresses joint admission control and per-user equipment (UE) bandwidth allocation to maximize weighted sum-rate in network slicing-enabled user-centric cell-free (CF) massive multiple-input multiple-output (mMIMO) systems when aggregate quality-of-service (QoS) demand may exceed available bandwidth. Specifically, we optimize bandwidth allocation while satisfying heterogeneous QoS requirements across enhanced mobile broadband (eMBB) and ultra-reliable low-latency communication (URLLC) slices in the uplink. The formulated problem is NP-hard, rendering global optimality computationally intractable. We decompose it into two sub-problems and solve them via computationally efficient heuristics within a sequential framework. We propose (i) a hierarchical admission control scheme that selectively admits UEs under bandwidth scarcity, prioritizing URLLC to ensure latency-sensitive QoS compliance, and (ii) an iterative gradient-based bandwidth allocation scheme that transfers bandwidth across slices guided by marginal utility and reallocates resources within slices. Simulation results demonstrate that the proposed scheme achieves near-optimal performance, deviating from an interior point solver-based benchmark by at most 2.2% in weighted sum-rate while reducing runtime by 99.7%, thereby enabling practical real-time deployment. Compared to a baseline round-robin scheme without admission control, the proposed approach achieves up to 1085% and 7% higher success rates for eMBB and URLLC slices, respectively, by intentionally sacrificing sum-rate to guarantee QoS. Sensitivity analysis further reveals that the proposed solution adapts effectively to diverse eMBB/URLLC traffic compositions, maintaining 47–51% eMBB and 93–94% URLLC success rates across varying load scenarios, confirming its robustness for resource-constrained large-scale deployments.
@@ I Introduction
The next-generation wireless systems aim to support diverse applications with stringent quality-of-service (QoS) demands, primarily represented by enhanced mobile broadband (eMBB) and ultra-reliable low-latency communication (URLLC) services [10]. These services exhibit fundamentally different characteristics: URLLC applications such as industrial automation and autonomous vehicles require strict latency bounds (e.g., 1 ms) and ultra-high reliability (99.999%) [1], where delays can cause critical failures, whereas eMBB traffic is elastic and delay-tolerant, allowing scheduling flexibility [14]. Furthermore, URLLC packets arrive sporadically with stringent deadlines, while eMBB demands sustained throughput over longer timescales [14]. Consequently, URLLC-first resource allocation ensures latency-critical traffic receives immediate access while eMBB user equipments (UEs) tolerate brief interruptions, enabling efficient coexistence under resource scarcity. Network slicing (NS) facilitates such service differentiation by partitioning physical infrastructure into logical slices optimized for distinct QoS objectives [2]. To realize these benefits, cell-free (CF) massive multiple-input multiple-output (mMIMO) has emerged as a potential sixth generation (6G) paradigm [11], where distributed access points (APs) cooperatively serve UEs, offering superior spectral efficiency (SE), uniform coverage, and enhanced reliability through user-centric (UC) connectivity [6]. However, combining NS with CF mMIMO introduces significant challenges in managing heterogeneous QoS requirements, particularly when prioritizing latency-sensitive URLLC traffic over eMBB services under limited spectrum.
@@
Building on this insight, we propose a joint admission control and bandwidth allocation framework with URLLC prioritization for NS-enabled uplink (UL) CF mMIMO systems under bandwidth scarcity where aggregate QoS demand may exceed available capacity. The framework operates in the presence of pilot contamination (PC) and ensures higher success rates for delay-sensitive URLLC UEs through selective admission and priority-aware resource allocation while maintaining system efficiency. To the best of our knowledge, joint priority-based admission control and bandwidth allocation in UC CF mMIMO systems have not yet been investigated in the literature. The main contributions are as follows:
$\bullet$ We formulate a joint optimization problem for admission control and per-UE bandwidth allocation to maximize weighted sum-rate while satisfying heterogeneous QoS constraints across eMBB and URLLC slices in NS-enabled UL CF mMIMO systems under bandwidth scarcity where aggregate QoS demand may exceed available capacity; owing to the NP-hard nature of the problem, we propose a sequentially optimization approach that decomposes it into two sub-problems and solves them in cascade.
@@ II-A Channel Propagation and Estimation
Each coherence interval of $\tau_{c}$ symbols allocates $\tau_{p}$ symbols for pilot transmission and the remainder for data. All UEs simultaneously transmit pilot sequences $\bm{\psi}_{k}\in\mathbb{C}^{\tau_{p}\times 1}$ with $\|\bm{\psi}_{k}\|^{2}=1$ assigned from a $\tau_{p}$ orthogonal sequences. Since $K\gg\tau_{p}$ in realistic deployments, pilot reuse introduces PC. We adopt minimum mean square error (MMSE) estimation to obtain channel estimates $\hat{\mathbf{g}}_{k,m}$, whose quality is quantified by $\gamma_{k,m}\triangleq\mathbb{E}\left[\|\hat{\mathbf{g}}_{k,m}\|^{2}\right]=\sqrt{\tau_{p}\rho_{p}\,\eta_{k}^{p}}\,\beta_{k,m}\,c_{k,m}$, where $\rho_{p}$ is the normalized pilot signal-to-noise ratio (SNR), $\eta_{k}^{p}\in(0,1]$ is UE $k$’s pilot power control coefficient [11], and
@@ II-B Uplink Data Transmission and Latency Characterization
Each UE $k\in\mathcal{K}$ transmits data symbols to APs in $\mathcal{V}_{k}$ [^1] Every AP captures signals from all UEs, yet it only processes the transmissions of its designated UEs as determined by the UE–AP association policy. during UL data transmission which forwards received signals to the CPU for centralized joint decoding using $\hat{\mathbf{g}}_{k,m}$. For eMBB services, Shannon capacity provides an accurate rate model, whereas URLLC short packets require finite blocklength analysis [17]. The achievable rate for both service classes is presented in (4) at the top of the next page, where $\theta$ is target decoding error probability, ${Q}^{-1}(\cdot)$ is the inverse Gaussian Q-function, and $V_{k}=1-(1+\text{SINR}_{k})^{-2}$ captures channel dispersion. Following [7], the signal-to-interference-plus-noise ratio (SINR) for UE $k$ is shown in (5), where $\rho_{d}$ is normalized UL data SNR and $\eta_{k}^{d}\in(0,1]$ is the data power control coefficient.
@@
To model URLLC end-to-end latency, we employ an $M/M/1$ queueing framework [15]. For UE $k\in\mathcal{K}_{\mathrm{URLLC}}$, packets arrive at rate $\lambda_{k}$ (packets/s) and are served at rate $\mu_{k}=R_{k}/L_{k}$ (packet size in bytes). Under stability condition $\mu_{k}>\lambda_{k}$ [9], the total delay is $D_{k}=\frac{1}{\mu_{k}-\lambda_{k}}=\frac{1}{\frac{R_{k}}{L_{k}}-\lambda_{k}}$, combining queueing and transmission delays. To satisfy URLLC latency requirements, we enforce $D_{k}\leq D_{k}^{\mathrm{max}}$, which translates to the minimum rate condition $\frac{R_{k}}{L_{k}}\geq\lambda_{k}+\frac{1}{D_{k}^{\mathrm{max}}},\forall k\in\mathcal{K}_{\mathrm{URLLC}}$, ensuring delays remain within bounds for mission-critical service delivery.
@@ III Problem Formulation
Problem $\mathcal{P}_{0}$ is a mixed-integer bilinear program, where binary $\bm{\alpha}$ variable induces combinatorial complexity of $2^{K}$, and product terms $\alpha_{k}b_{k}$ render (6a) and (6d) non-convex. However, encoding strict lexicographic priority via constraints would require $\mathcal{O}(|\mathcal{K}_{\text{URLLC}}|\times|\mathcal{K}_{\text{eMBB}}|)$ binary conditions, further exacerbating complexity.
@@ IV-A2 Computational Complexity
$B_{\text{residual}}\leftarrow B-B_{\text{URLLC}}^{\text{used}}$
@@ IV-B1 Proposed Bandwidth Allocation Scheme
The iterative loop (Lines 22) finalizes the bandwidth allocation. Algorithm 2 uses another subroutine for calculating marginal utility, $\mu_{s},\ s\in\mathcal{S}$ in Line 2. From these $\mu_{s}$ values, the algorithm identifies donor (minimum $\mu_{s}$) and receiver (maximum $\mu_{s}$) slices (Line 2). The algorithm terminates if marginal utilities are balanced ($\mu_{s_{\text{rcv}}}\leq 1.05\mu_{s_{\text{donor}}}$) in Line 2. Otherwise, bandwidth $\Delta B$ is transferred from donor to receiver slice (Line 2), where $\Delta B$ is the average per-UE bandwidth in the donor slice, capped by the excess above its minimum requirement. Termination also occurs if $\Delta B$ is below a small threshold $\epsilon_{\text{transfer}}$ (Line 2). After updating slice allocations (Line 2), per-UE bandwidth is again recomputed with QA subroutine in Line 2, and Line 2 evaluates the new objective $\tilde{f}=\sum_{k\in\mathcal{K}_{\text{adm}}}w_{k}\tilde{b}_{k}\text{SE}_{k}$. The transfer is accepted if $\tilde{f}$ exceeds the previous objective (Lines 22); otherwise, it is rejected (Line 2). However, early stopping can also occur after $T_{\text{patience}}$ consecutive rejections (Line 2). Algorithm 2 employs two subroutines for per‑UE bandwidth calculation in Lines 2 and 2, and marginal‑utility computation in Line 2, described next.
@@
Algorithm 2 initialization requires $\mathcal{O}(S+K)$ for surplus distribution and initial allocation. Each iteration incurs $\mathcal{O}(SK)$ for marginal utility computation (Line 2), $\mathcal{O}(S)$ for donor-receiver identification, and $\mathcal{O}(K)$ for reallocation and evaluation (Lines 2 and 2), yielding a per-iteration complexity $\mathcal{O}(SK)$. Over at most $T_{\max}$ iterations, the overall computational complexity of Algorithm 2 is $\mathcal{O}(T_{\max}SK)$.
@@ V-1 Average Weighted Sum-Rate Performance
Fig. 1 compares the average weighted sum-rate performance among three schemes: the proposed scheme (‘Proposed’), a ‘CVX’ scheme employing CVX, a disciplined convex programming solver-based bandwidth allocation, and a ‘Baseline’ scheme using round-robin bandwidth allocation without admission control. Both the CVX and Proposed schemes utilize the proposed priority-based admission control (Algorithm 1) and the sequential optimization technique to solve $\mathcal{P}_{0}$.
@@ V-2 Average Success Rate Performance
Figs. 2 and 3 illustrate the average success rate, defined as the fraction of UEs satisfying their QoS constraints (minimum-rate for eMBB, delay for URLLC). Importantly, Proposed and CVX achieve identical success rates, confirming that Algorithm 2 preserves QoS feasibility while matching the solution quality nearly identical to the CVX optimum. Compared to Baseline, Proposed outperforms it by 1085% for eMBB and 7% for URLLC. Note that URLLC success rates consistently exceed eMBB due to smaller bandwidth demands per packet and URLLC-prioritized admission control. As $K$ increases, success rates degrade for all schemes due to resource contention, however, Proposed maintains substantially higher QoS compliance. This advantage stems from the synergy between Algorithm 1 and Algorithm 2, which jointly allocate bandwidth based on slice-specific weights and SE, directing resources toward high-priority UEs capable of meeting stringent constraints.
@@ V-4 Sensitivity Analysis of the Proposed Solution
Fig. 5 examines the sensitivity of the Proposed scheme under varying slice compositions, where Low, Medium, and High loads correspond to 70%/30%, 50%/50%, and 30%/70% eMBB/URLLC UE ratios, respectively. Notably, Proposed maintains near-optimal solution quality, deviating from CVX by at most 2.0% across all loads, validating the effectiveness of Algorithm 2. Compared to Baseline, the Proposed scheme intentionally sacrifices average weighted sum-rate by 7.3%, 15.7%, and 35.5% at Low, Medium, and High loads, respectively, to guarantee QoS compliance. However, this modest sum-rate penalty yields substantial QoS benefits. For eMBB UEs, Proposed achieves 1371%, 365%, and 98% higher success rates than Baseline at Low, Medium, and High loads, respectively, while maintaining stable performance (47-51%) across all loads. For URLLC UEs, Proposed outperforms Baseline by 7.0%, 9.5%, and 13.3%, sustaining 93-94% success rates even under High load with 70% URLLC demand. This behavior demonstrates that the URLLC-prioritized admission control (Algorithm 1), complemented by Algorithm 2, effectively adapts to diverse slice mixes by directing resources toward high-priority, latency-sensitive UEs, thereby ensuring robust QoS compliance at the expense of raw sum-rate.
Figure 5: Sensitivity analysis of the different schemes under varying load with $K=100$ UE and $M=100$ APs.